# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v1

@@ 3.3 Training Calibration Loss
We then calibrate the FEU model of the epistemic uncertainty against predictive accuracy. Specifically for the MoE top-k routers, we design the following calibration loss for training:
@@
where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$. As shown in Figure 1, this term effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness. Therefore, this calibrated training loss ensures that the router’s certainty aligns more closely with the actual confidence the model should have given the available data and its knowledge state. Thus, it effectively calibrates the router’s decisions in line with epistemic uncertainty.
Load balancing is a common technique to ensure even exploitation across experts with the MoE architecture (Fedus et al., 2022). We follow the load balancing loss $\mathcal{L}_{b}$ proposed by (Li et al., 2024) and define our loss function as:
$\mathcal{L}=\mathbf{CE}+\alpha\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (9)
where $\mathbf{CE}$ represents cross entropy loss, $\alpha$ and $\beta$ are the hyperparameters of two auxiliary terms.
@@ 4.1 Parameter-Efficient Fine Tuning for LLMs
Large Language Models (Brown et al., 2020; Chowdhery et al., 2022; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks. Building on these advances, instruction fine-tuning(Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems(Wu et al., 2023b; Achiam et al., 2023b). However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT)(Mangrulkar et al., 2022), model distillation(Liu et al., 2023; Xiao et al., 2023), quantization(Frantar et al., 2022; Xiao et al., 2022b), and pruning(Frantar and Alistarh, 2023; Ma et al., 2023). Among these, LoRA(Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference. For example, VeRA(Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara(Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings. Tied-Lora(Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters. AdaLoRA(Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA(Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
@@ 4.2 Mixture of Experts
The Mixture-of-Experts concept(Jacobs et al., 1991), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data. Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies. For example, LLaVA-MoLE(Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines. Other MoE-based approaches include MoRAL(Yang et al., 2024b), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE(Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting. PESC(Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational and GPU memory requirements. MoE-LoRA(Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE(Gou et al., 2023) activates task-specific model parameters based on instruction clusters. MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
@@ 4.3 Uncertainty Quantification in LLMs
Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks(Gal and Ghahramani, 2015, 2016b; Malinin and Gales, 2018; Ovadia et al., 2019; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023). This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research(Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023). LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs(Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs(Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
Existing works have investigated the application of ensembling in fine-tuning LLMs for uncertainty quantification. Gleave and Irving (2022); Sun et al. (2022) focus on uncertainty estimation in full model fine-tuning, while this approach inherently incurs significant memory overhead. Wang et al. (2023); Zhai et al. (2023b); Balabanov and Linander (2024) explore the use of LoRA ensembles for uncertainty estimation in LLMs. Yang et al. (2024a) applies a post-hoc Laplace approximation Mackay (1992) to model LoRA parameters after fine-tuning. BatchEnsemble(Wen et al., 2020), introduces component-specific rank-1 matrices as multiplicative modifications to a base model. Though this method has been applied to LLMs, it has been used in the pre-training phase rather than fine-tuning(Tran et al., 2022). None of these methods provide calibrations on epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
@@ 5.1 Datasets
We include $5$ multiple-choice question answering benchmarks to evaluate UQ4CT: OpenBookQA (OBQA, Mihaylov et al. (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science. We also use computer science, law, medication and engineering subsets from MMLU dataset(Hendrycks et al., 2020) to evaluate performance under distribution shift. We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance.
Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across $4$ common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
@@ 5.2 Experiment Setup
  • Monte Carlo (MC) Dropout(Gal and Ghahramani, 2016a) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
@@
  • Laplace-LoRA (LA)(Yang et al., 2024a) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
@@
We measure the prediction accuracy on the validation set for all $5$ tasks. For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al. (2017)) with bin size of $15$, which measures the alignment between predicted probabilities and empirical accuracy. We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications. Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts. Metric details are provided in Appendix A.2.
@@ 5.3 Results
Table 2: Performance comparison of different methods fine-tuned on OBQA dataset with LlaMA2-7B across $2$ smaller distribution shift (DS) tasks and $4$ larger distribution shift tasks. UQ4CT shows significant ECE improvements while maintaining high accuracy.
@@ 5.3.1 In-distribution Performance
The most significant performance improvement is observed in the reduction of Expected Calibration Error (ECE). UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks. Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration. This improvement is critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
@@ 5.3.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios. Similar to the dataset setup in (Yang et al., 2024a), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA. For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals. This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task. Details of the ensemble is provided in Appendix. A.3.
The distribution shift evaluations are provided in Table 2. UQ4CT provides significant improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios. For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario. For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models. Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks. This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
@@ 5.4.1 Incremental Weighting on Calibration Term
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintains similar accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA

### v2

@@ 3.4 Training Calibration Loss
The FEU model of the epistemic uncertainty enables us to calibrate the parameter mixtures against predictive accuracy for better aligned output distributions. Specifically for the MoE top-k routers, we design the following calibration loss for training:
@@
where the first term is an indicator function of whether the model prediction matches the ground truth $y^{*}$ given the prompt $x$. We employ a one-hot definition of the ground truth confidence. When the prediction from current mixture model matches the ground truth, the ground truth confidence is $1$. Otherwise, when the predictions do not match, the ground truth confidence is 00.
3.5 Calibrating the Mixture
Our proposed calibration loss $\mathcal{L}_{\mathrm{cal}}$ improves predictive reliability by adaptively balancing expert exploitation and exploration with functional-level epistemic uncertainty. As shown in Figure 1, $\mathcal{L}_{\mathrm{cal}}$ effectively promotes expert exploitation for correct predictions and expert exploration for incorrect predictions by directly calibrating the functional level epistemic uncertainty to align with the predictive correctness.
Ideally, when the $K$ LoRA experts together capture all the functional relationships across the data distribution with cross entropy during fine-tuning, our proposed loss $\mathcal{L}_{cal}$ also finds proper mixture of LoRA experts conditioned on the input $x$ by conditionally promoting expert exploitation and exploration. This allows the model to select correct functional relationships regarding $x$ to generate an output that better matches the data distribution, which grants calibrated uncertainty estimations.
Load balancing is a common regularization to ensure even utilization across experts with the MoE architecture (Fedus et al., 2022). We follow the load balancing loss $\mathcal{L}_{b}$ proposed by Li et al. (2024) and define our loss function as:
$\mathcal{L}=\mathbf{CE}+\gamma\cdot\mathcal{L}_{b}+\beta\cdot\mathcal{L}_{\mathrm{cal}},$ (14)
where $\mathbf{CE}$ represents cross entropy loss, $\gamma$ and $\beta$ are the hyperparameters of two auxiliary terms. Details about $\mathcal{L}_{b}$ can be found in Appendix A.1.
@@ 4.1 Parameter-Efficient Fine Tuning for LLMs
Large Language Models (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023b, c) have shown impressive abilities in handling various natural language processing tasks. Building on these advances, instruction fine-tuning (Chung et al., 2022; Iyer et al., 2022; Zheng et al., 2024) has enhanced LLMs’ capacity to comprehend and follow human instructions, forming the core of modern conversational AI systems (Wu et al., 2023b; Achiam et al., 2023b). However, as LLMs increase in size, the fine-tuning process demands much more time and memory.
To address these challenges, several strategies have been proposed, including parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022), model distillation (Liu et al., 2023; Xiao et al., 2023), quantization (Frantar et al., 2022; Xiao et al., 2022b), and pruning (Ma et al., 2023). Among these, LoRA (Hu et al., 2021b), which leverages low-rank matrix decomposition of linear layer weights, is a widely adopted PEFT technique that boosts model performance without adding computational costs during inference. For example, VeRA (Kopiczko et al., 2023) introduces learnable scaling vectors to modify shared pairs of frozen random matrices across layers, while FedPara (Hyeon-Woo et al., 2021) focuses on low-rank Hadamard products for federated learning settings. Tied-Lora (Renduchintala et al., 2023) applies weight tying to further minimize the number of trainable parameters. AdaLoRA (Zhang et al., 2023) uses Singular Value Decomposition (SVD) to prune less important singular values for efficient updates, and DoRA (Liu et al., 2024) separates pre-trained weights into magnitude and direction components, applying LoRA to update the directional component during fine-tuning, thus reducing the number of parameters to be trained.
@@ 4.2 Mixture of Experts
The Mixture-of-Experts concept (Jacobs et al., 1991; Wang et al., 2020), introduced as early as 1991, presented a novel supervised learning framework where multiple networks (experts) specialize in handling distinct subsets of training data. Modern MoE variants adapt this by modifying the traditional feed-forward sub-layer within transformer blocks, incorporating sparsely activated LoRA experts, which allows for significant expansion in model width without a proportional increase in computational overhead.
Different MoE architectures have since emerged, distinguished by their expert sampling and routing strategies. For example, LLaVA-MoLE (Chen et al., 2024) improves token routing to domain-specific experts within transformer layers, reducing data conflicts and consistently outperforming standard LoRA baselines. Other MoE-based approaches include MoRAL (Yang et al., 2024a), which focuses on efficiently adapting LLMs to new domains and tasks for lifelong learning, and LoRAMoE (Dou et al., 2024), which incorporates LoRAs via a router network to mitigate the issue of world knowledge forgetting. PESC (Wu et al., 2024b) transforms dense models into sparse ones through an MoE structure, lowering computational costs. MoE-LoRA (Luo et al., 2024) introduces a new parameter-efficient MoE method using Layer-wise Expert Allocation (MoLA) for transformer models, while MoCLE (Gou et al., 2023) activates task-specific model parameters based on instruction clusters. MixLoRA (Li et al., 2024) implements a high-throughput framework for LoRA MoE training and inference process, constructing LoRAs as stochastic experts to reduce computational overhead while expanding model capacity.
@@ 4.3 Uncertainty Quantification in LLMs
Uncertainty quantification has garnered substantial attention in various tasks and domains within neural networks (Gal and Ghahramani, 2015, 2016a; Malinin and Gales, 2018; Ovadia et al., 2019; Dusenberry et al., 2020; Malinin et al., 2021; Lin et al., 2022; Kuhn et al., 2023; Lin et al., 2023). This focus extends to LLMs, where the precise quantification of prediction uncertainty has become a critical area of research (Xiao et al., 2022c; Lin et al., 2022; Mielke et al., 2022; Chen and Mueller, 2023; Duan et al., 2023; Huang et al., 2023). LLMs, particularly in generative tasks, pose unique challenges, especially when it comes to measuring the uncertainty of their outputs (Liu et al., 2019; Malinin and Gales, 2021; Kuhn et al., 2023; Lin et al., 2023). The distinction between aleatoric and epistemic uncertainty was recently examined in the context of LLMs (Hou et al., 2023), though this was approached by ensembling model inputs rather than model instances and did not address fine-tuning tasks specifically.
Established uncertainty quantification methods have been studied in conjunction with the LoRA structure. Monte-Carlo dropout (Gal and Ghahramani, 2016b) interprets dropout in neural networks as approximate Bayesian inference in deep Gaussian processes, allowing uncertainty estimates to be obtained from existing LoRA adapters without modifying them. Checkpoint ensemble (Chen et al., 2017) utilizes predictions from multiple LoRA checkpoints saved during a single fine-tuning process to calibrate uncertainty. Deep ensemble (Lakshminarayanan et al., 2017; Wang et al., 2023; Zhai et al., 2023) combines the predictions from multiple LoRA adapters for better uncertainty calibration. Laplace-LoRA (Yang et al., 2024b) applies Bayesian inference via Laplace approximation to the LoRA parameters after fine-tuning, resulting in improved calibration and uncertainty estimates. However, none of these methods provide calibrations regarding epistemic uncertainty, which is crucial to mitigate overconfidence in the fine-tuning stage given the sparse dataset.
@@ 5.1 Datasets
We include $5$ multiple-choice question answering benchmarks: OpenBookQA (OBQA, Mihaylov et al. (2018)), ARC-Easy (ARC-E) and ARC-Challenge (ARC-C) from AI2 Reasoning Challenge (Clark et al., 2018), BOOLQ (Clark et al., 2019) and ClimateQA, an expert-annotated domain specific benchmark for climate science. We also use computer science, law, medication and engineering subsets from MMLU dataset (Hendrycks et al., 2020) to evaluate performance under distribution shift. We fine-tune on the publicly available training split and test on the validation split from these benchmarks to evaluate model performance. We report the average model performance over $3$ random runs and the standard deviations in the subscript.
Table 1: Performance comparison of different methods fine-tuned with LlaMA2-7B across four common sense reasoning tasks and a domain-specific task. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@ 5.2 Experiment Setup
  • Monte Carlo (MC) Dropout (Gal and Ghahramani, 2016b) keeps dropout on at both training and testing time. By performing multiple forward passes, MC dropout randomly shuts down a portion of model nodes, producing ensemble-alike predictions. To combine LoRA fine-tuning with MC dropout, we apply dropout on the input of the LoRA adapter, following the implementation of Mangrulkar et al. (2022).
@@
  • Laplace-LoRA (LA) (Yang et al., 2024b) applies a post-hoc Laplace approximation on fine-tuned LoRA parameters for better uncertainty estimation.
@@
We measure the prediction accuracy on the validation set for all $5$ tasks. For uncertainty calibration, we incorporate expected calibration error (ECE, Guo et al. (2017)) with $15$ bins, which measures the alignment between predicted probabilities and empirical accuracy. We also investigate model performance under distribution shift to ensure the model has predictable behavior when given data from other domains as this is a crucial component for real-world applications. Specifically, we test models fine-tuned on OBQA dataset with $4$ domain-specific MMLU subtask ensembles focusing on different professionalities and ARC-E/C datasets to approximate larger and smaller distribution shifts. Metric details are given in Appendix A.6.
@@ 5.3 Results
Table 2: Performance comparison of different methods fine-tuned on the OBQA dataset with LlaMA2-7B across two smaller distribution shift (DS) tasks and four larger distribution shift tasks. UQ4CT shows substantial ECE improvements while maintaining high accuracy.
@@ 5.3.1 In-distribution Performance
The most substantial performance improvement is observed in the reduction of Expected Calibration Error (ECE). UQ4CT consistently outperforms other methods, reducing ECE by more than $25\%$ on average across the evaluated benchmarks. Unlike other methods where the ECE performance is worsened on the more challenging ARC-C benchmark, UQ4CT achieves an ECE score of $6.1$, showcasing the effectiveness of the calibration.
In addition to experiments on LLaMA-2-7B in the main text, we also present additional experiments on fine-tuning Mistral-7B in Appendix A.2 for more comprehensive evaluation of our method. For both LLaMA-2-7B and Mistral-7B models, UQ4CT consistently shows substantial improvements in uncertainty calibration. The improvements are critical in applications where the model’s confidence must align with its predictive accuracy given limited data, particularly in safety-critical and domain-specific tasks.
@@ 5.3.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios. Similar to the dataset setup in (Yang et al., 2024b), we use ARC-C and ARC-E dataset to simulate smaller distribution shift because the ARC dataset has similar domain focus on general science reasoning, but is generally more challenging and covers a broader range of scientific topics than OBQA. For larger distribution shift, we ensemble the domain-specific MMLU subtasks into $4$ benchmarks focusing on different professionalities: Computer Science (CS), Engineering (Eng), Law and Health. These tasks have very broad coverage of the domain task at various knowledge levels ranging from elementary school to professionals. This domain-specificity demonstrates larger distribution shift from OBQA, which is a general common sense reasoning task. Details of the ensemble is provided in Appendix. A.7.
The distribution shift evaluations are provided in Table 2. UQ4CT provides substantial improvements in terms of ECE while maintains similar accuracy for both smaller and larger distribution shift scenarios. For smaller distribution shifts, UQ4CT shows comparable ECE performance as the in-distribution scenario. For the more challenging larger distribution shifts, UQ4CT still achieves the best ECE performance among all baseline models. Note that UQ4CT also achieves competitive prediction accuracy across all domain-specific tasks. This empirically shows that our proposed alignment of the functional epistemic uncertainty with predictive correctness improves generalizability and mitigating the overconfidence problem on the fine-tuned model.
Table 3: Performance comparison of UQ4CT with and without incremental weighting. Incremental weighting has worse ECE performance while maintaining similar accuracy.
Metrics Methods BoolQ ARC-E ARC-C OBQA ClimateQA
@@ 5.4 Ablation Study
In this section, we conduct ablation studies to investigate the effectiveness of our designed calibration loss, $\mathcal{L}_{cal}$. We first evaluate the incremental weighting performance of the calibration term, which investigates the effectiveness of $\mathcal{L}_{cal}$ at the early stage of fine-tuning. Then, we perform a sensitivity test, where we explore the overall performance impact of $\mathcal{L}_{cal}$. We also conduct an ablation study on the impact of active LoRA experts in Appendix A.4.
@@ 5.4.1 Incremental Weighting on Calibration Term
Table 4: Performance of UQ4CT with varying $\beta$ values on the OBQA dataset. Prediction accuracy and uncertainty calibration improve with increasing $\beta$, highlighting the effectiveness of $\mathcal{L}_{\mathrm{cal}}$.
$\beta$ ACC $\uparrow$ ECE $\downarrow$
0 $75.5_{2.91}$ $12.9_{1.99}$
0.2 $76.0_{0.6}$ $7.47_{0.78}$
0.5 $75.9_{0.31}$ $7.82_{0.93}$
0.8 $76.6_{0.4}$ $5.94_{1.16}$
1 $\mathbf{77.3_{1.36}}$ $\mathbf{5.0_{1.15}}$
5.4.2 Sensitivity Test on Calibration Term
To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values of $0,0.2,0.5,0.8$ and $1$, where $\beta=0$ resembles the original MixLoRA method.
Results in Table 4 demonstrate the effectiveness of the calibration loss. When $\beta=0$, the model is optimized without calibration on parameter mixtures, resulting in high ECE value. Even with small $\beta=0.2$ or $\beta=0.5$, the ECE scores drastically improved compared to no calibration setting. Finally, when $\beta=1$, the calibration term effectively optimizes the conditional parameter mixtures to generate outputs that fit data distribution well, resulting in lower ECE scores and higher accuracies.